# Chain-of-Authorization: Internalizing Authorization into Large Language Models via Reasoning Trajectories

### 24 Mar 2026 v1

@@ 2.2.2 Learning Paradigm: Internalizing Authorization via Supervised Fine-tuning
Based on the synthetic data described above, we employ a unified supervised fine-tuning framework. By minimizing the negative log-likelihood loss under the same sequence prediction objective, we strongly bind the authorization decision to the response generation:
@@ 3 Evaluation
In this section, by analyzing and presenting experimental results of Chain-of-Authorization (CoA), we discuss whether CoA can endow Large Language Models (LLMs) with robust intrinsic access control capabilities without compromising their performance on downstream tasks.
@@ 3.1 Model Utility on Authorized States
In this section, we evaluate the performance of the proposed Chain of Authorization (CoA) framework under authorization constraints across three different model architectures (Qwen3-1.7B, Llama3.1-8B-instruct, and Mistral-7B-v0.3). The evaluation benchmarks are divided into three categories: internal parametric knowledge tasks, including WMDP [li2024wmdp] and MMLU [hendryckstest2021mmlu]; external context tasks, including SQuAD [rajpurkar-etal-2016-squad] and CovidQA [covidqa, moller-etal-2020-covidqa]; and tool-calling tasks, represented by Mobile-Actions [google_mobile_actions]. We use accuracy to evaluate each model’s utility and policy adherence.
We compare CoA against several baselines, including the Base model, direct supervised fine-tuning (SFT), extra classification model (SFT+Extra), PermissionLM (PermLM), and sudoLM. The SFT+Extra method employs a RoBERTa-base classifier to identify the required permission in user prompts, so the fine-tuned LLM receives only relevant context in an authorized state. PermLM trains submodels for WMDP and MMLU and employs a context-filtering strategy in SQuAD and CovidQA to ensure LLMs receive contexts only from authorized documents.
@@ 3.2 Authorization under Different Permissions
Figure 3: Accuracy across five datasets under mismatch and public authorization states, where each circle represents the average accuracy of a method on a group of specific datasets across three backbone models, with the line mapping to its variance.
Figure 3 illustrates the accuracy of different methods in the unauthorized state. Since high accuracy in the unauthorized scenario usually indicates that unauthorized information has been successfully utilized, the ideal authorized method should result in the ”Optimal Zone” in the lower-left corner, that is, maintaining extremely low accuracy in both the public and mismatch states. Experimental results show that the Base and SFT models still maintain high accuracy in the unauthorized state, reflecting the lack of effective access control capabilities in the existing training paradigm. Although slightly lower than SFT in the public state, PermissionLM still retains high accuracy, failing to completely block information leakage. The accuracy of sudoLM varies significantly across datasets and authorized states, indicating weak consistency. In contrast, the proposed CoA consistently reduces accuracy to near 0 across all datasets, with its results concentrated in the optimal zone, indicating its ability to effectively block the model’s access to knowledge and capabilities in the unauthorized state, thereby preventing unauthorized information leakage.
Figure 4: Refusal rate across five datasets under mismatched and public authorization state, where each circle represents the average refusal rate of a method on a group of specific dataset across three backbone models, with the line mapping to its variance.
Figure 4 illustrates the rejection rate of different methods in different unauthorized states. An ideal model should be located in the ”Optimal Zone” in the upper-right corner, maintaining a high rejection rate in both public and mismatched unauthorized states. As shown in Figure 4, the base model and SFT models lack basic authorization awareness, with an almost zero rejection rate across all datasets. While PermissionLM restricts information access through structural isolation, it also lacks an explicit rejection mechanism. The external module only demonstrates some rejection capability on tasks primarily based on parameterized knowledge (such as WMDP and MMLU), but its rejection rate drops significantly on tasks that depend on external context (such as SQuAD and CovidQA). Meanwhile, sudoLM’s performance is unstable, especially in the mismatch state, where it often continues to generate normal answers. In contrast, our proposed CoA consistently falls within the optimal zone on all models and datasets, achieving a rejection rate of nearly 100% in both the public and mismatch unauthorized states, demonstrating its ability to reliably identify fine-grained authorization states.
@@ 3.3 Robustness Against Adversarial Prompts
Table 2: Attack Success Rate (%) under the mismatch authorization state on WMDP. Bold and italicized values indicate the best and second-best results, respectively.
@@
None 100.00 0.14 78.64 0.14 98.50 0.00
Prefix Injection 100.00 0.14 79.18 0.14 0.95 0.27
Style Injection 100.00 0.14 51.43 0.14 98.10 0.14
Misrepresentation 99.86 0.14 57.28 0.00 97.28 0.14
Logical Appeal 100.00 0.00 72.38 0.14 98.24 0.00
Authority Endorsement 100.00 0.00 66.12 0.00 97.42 0.00
Expert Endorsement 100.00 0.27 48.03 0.27 97.15 0.00
Evidence Persuasion 100.00 0.00 69.52 0.00 98.51 0.14
PAIR 6.39 0.00 0.27 0.00 12.52 0.00
As shown in Table 2, under the mismatch authorization state, CoA exhibited strong defensive consistency, obtaining the optimal or near-optimal ASR in most scenarios. Particularly in the style injection scenario, where the DPO baseline has an obvious vulnerability, CoA successfully reduces the ASR of the Llama-3-8B-Instruct from $52.52\%$ to $0.14\%$. This cross-model robustness demonstrates that CoA can effectively filter out semantic perturbations intended to circumvent authorization policies by embedding authorization decisions into the reasoning trajectory.
Table 3: Attack Success Rate (%) under the public authorization state on WMDP. Bold and italicized values indicate the best and second-best results, respectively.
ASR($\downarrow$) Qwen3-1.7B Llama3-8B-Instruct Mistral-7B-v0.3
SudoLM CoA SudoLM CoA SudoLM CoA
None 3.67 6.54 9.25 0.00 17.82 0.00
Prefix Injection 3.27 6.80 0.00 0.00 0.00 0.00
Style Injection 99.86 9.12 7.89 0.14 66.12 0.00
Misrepresentation 0.00 5.58 4.90 0.00 1.77 0.00
Logical Appeal 0.00 5.44 3.68 0.00 0.69 0.00
Authority Endorsement 0.00 5.31 1.50 0.00 0.41 0.00
Expert Endorsement 0.00 4.49 2.32 0.00 0.82 0.00
Evidence Persuasion 0.28 5.44 1.91 0.00 0.14 0.00
PAIR 0.82 0.68 0.27 0.27 2.99 0.00
Table 3, on the other hand, demonstrates the robustness results under the public authorization state, where CoA maintained a robust defense boundary. Despite facing persuasive attacks driven by GPT-4o, CoA repeatedly achieved an ASR of $0.00\%$ on both the Llama-3.1-8B-Instruct and Mistral-7B-v0.3. In contrast, sudoLM showed significant fluctuations with style injection, with its ASR rising to $66.12\%$ on Mistral-7B-v0.3, while CoA remained robust at $0.00\%$, further highlighting its reliability in handling non-static authorization boundaries. In summary, experiments under two authorization states jointly verify that CoA can effectively resist diverse malicious attacks across models of different sizes.
@@ 3.4 Visualization: What does CoA do?
To delve deeper into how the CoA mechanism reshapes the model’s internal representation space, we conducted a visualization analysis of the fine-tuned Qwen3-1.7B model on the WMDP dataset, comparing it with its original version. First, we constructed three scenarios with clear real-world authorization implications for each sample in the test set:
@@
  • Public: We add the ”¡—public—¿” label to all samples to simulate a visitor’s access scenario.
Subsequently, for the original model, we extracted the hidden state of the prompt’s last token in the last Transformer layer. For the fine-tuned model, we extracted features at two positions: one is the last token of the prompt, at which point the model has absorbed complete semantic and permission input information; the other is the last token before the authorization reasoning process ends and the final answer is generated, containing the authorization decision information after reasoning. Finally, we first use PCA to reduce the dimension to $50$ to suppress noise and improve computational efficiency [^1] We perform PCA first, following the official guidance of scikit-learn library https://scikit-learn.org/stable/modules/generated/sklearn.manifold.TSNE.html., and then use t-SNE to reduce the dimension to $2$ for visualization.
@@
The visualization results are shown in Figure 5. The original model (prompt end) shows the distribution of representations at the end of the prompt. Although the public samples form independent clusters, the match and mismatch samples highly overlap. This indicates that the original model cannot semantically distinguish the prompt under different authorization states. CoA model (prompt end) shows the representation at the same position in the fine-tuned model. Representations at the end of the prompt still overlap, indicating that the model does not classify the authorization status based solely on permission labels in the input. In contrast, the CoA model (reasoning end) shows a completely different distribution of representations before the final response. The three sample types form distinct, non-overlapping clusters. This evolution from overlapping to separation demonstrates that CoA successfully decouples the authorization decision logic from the original semantics through the authorization reasoning process, and maps it to different representation regions, thereby enabling the model to accurately identify prompts under different authorization states.
@@ 3.5 Causal Analysis of CoA Trajectories
Figure 6 illustrates the evolution of the reasoning trajectory of CoA under targeted intervention and its causal impact on the final answer. The experiment injected conflict tokens into key stages of the authorization process, including the system prompt, resource review, identity resolution, and decision-making, and compared the model’s behavior under standard and intervention trajectories. The results show that under the normal baseline without intervention, the model can stably complete the task without rejection. However, when any key stage exhibits logical inconsistency (e.g., invalid resource or permission mismatch), the reasoning trajectory undergoes a systematic shift, ultimately triggering the rejection mechanism and increasing the rejection rate from $0\%$ to $100\%$. This result indicates that the security of CoA does not stem from superficial prompt constraints, but is driven by the causal dependencies between logical states in the authorization process. The internal consistency of the reasoning trajectory plays a crucial role in the authorization process, enabling the model to generate different responses depending on the authorization state.
@@ 3.6 Ablation Study
To systematically evaluate the impact of hyperparameters on model performance, we conducted experiments on the WMDP dataset using the Llama3-8B-Instruct model with a batch size of $4$ and an epoch of $5$. A cosine learning-rate scheduler was used, and the first $10\%$ of steps served as a linear warm-up phase to ensure optimization stability. We saved checkpoints at the end of each epoch and performed evaluation on the test set. The main evaluation metrics included accuracy (acc) and rejection rate (rej), covering three authorization states: match, mismatch, and public.
@@ 5 Methods
In this section, we will describe the theoretical architecture and implementation path of Chain-of-Authorization. We first formally define the data access control problem in the context of Large Language Models (LLM), and then delve into the core mechanism design and the corresponding learning paradigm.
5.1 Problem Definition
In the standard autoregressive generative paradigm, the cognitive and reasoning processes of a large language model can be formalized as a conditional probability mapping from a high-dimensional input information space $\mathcal{X}$ to an output space $\mathcal{Y}$. Given a complete input sequence $X\in\mathcal{X}$, this sequence is typically constructed in semantic space by concatenating the user prompt $Q$, the external context $E$, and the set of available tools $G$. The model is parameterized by $\theta$, and its standard generative behavior can be represented as a conditional probability distribution over the vocabulary:
$\pi_{\theta}(Y|X)=\prod_{t=1}^{|Y|}P_{\theta}(y_{t}|X,y_{<t})$ (5)
Under this paradigm, the model treats all tokens in $X$ as equally accessible contextual backgrounds, lacking intrinsic awareness of information ownership and access boundaries.
To establish access boundaries for information flow, we introduce a global permission-label space $\mathcal{C}$. Every element in the information space $\mathcal{X}$ must be mapped to this permission space to establish its access threshold. Specifically, the permissions required to access sensitive internal knowledge, read external context $E$, and available tools $G$ to answer user prompt $Q$ are defined as subsets $C_{q},C_{e},C_{g}\subseteq\mathcal{C}$. The comprehensive set of permission conditions required to complete the current complex reasoning task constitutes the task requirement set $C_{req}=C_{q}\cup C_{e}\cup C_{g}$. Meanwhile, users are assigned identity credentials, defined as a set of user permissions $C_{u}\subseteq\mathcal{C}$. Therefore, the core logic of access control is abstracted into a policy evaluation function $\phi$, used to accurately calculate the inclusion relationship between user credentials and the overall task requirements:
$\phi(C_{req},C_{u})=\begin{cases}1,&\text{if }C_{req}\subseteq C_{u}\\ 0,&\text{otherwise}\end{cases}$ (6)
Under this formal framework, our goal is to construct a conditional probability distribution $\pi_{\theta}(Y|Q,E,G,C)$ that incorporates permission-aware capabilities. This distribution must be strictly subject to the mathematical constraints of the policy function $\phi$. We require the model to simultaneously satisfy the following utility and security constraints during the generation phase:
  • Utility constraint: When $\phi(\{C_{e}\cup C_{q}\cup C_{g}\},C_{u})=1$, the model output should approximate the original unrestricted distribution, i.e., $\pi_{\theta}(Y|Q,E,G,C)\approx\pi_{\theta}(Y|Q,E,G)$.
  • Security constraint: When $\phi(\{C_{e}\cup C_{q}\cup C_{g}\},C_{u})=0$, the model output distribution should be redirected to the rejection response distribution $\delta_{rej}$.
The fundamental obstacle to achieving the aforementioned conditional distribution lies in the structural decoupling between the model’s underlying cognitive mechanism and external policy verification. When $\phi(C_{req},C_{u})=0$, its underlying cause is heterogeneous: it may originate from an unauthorized attempt by the user ($C_{u}=\{c_{pub}\}$), or it may originate from a permission mismatch where the user holds some permission labels but not enough to cover specific confidential knowledge ($C_{req}\not\subseteq C_{u}$).
Traditional access control paradigms treat $\phi$ as an external post-hoc filter or a shallow prompt-level constraint, preventing the model from forming an internal causal linkage between the authorization state $C_{u}$ and the final response. As a result, under permission mismatch or adversarial context perturbations, the model is prone to feature-level ambiguity, leading to deviations toward unauthorized outputs rather than $\delta_{rej}(Y)$. In the CoA framework, leveraging the auto-regressive nature of LLMs, the LLM and policy function are unified as a joint distribution $\pi_{\theta}(Y,T\mid X,E,C)$, where the policy execution is instantiated as an explicit authorization inference trajectory $T$ (the chain-of-authorization), which serves as the causal premise for generating the response $Y$:
$\pi_{\theta}(Y,T|X,E,G,C)=\pi_{\theta}(T|X,E,G,C)\cdot\pi_{\theta}(Y|T;X,E,G,C)$ (7)
With this formalization, the chain-of-authorization $T$ determines the direction of the final response: when $\phi(\{C_{q},C_{e},C_{g}\},C_{u})=1$, $T$ generates a decision for the authorization state and guides the LLM to generate the expected response; conversely, if $T$ generates a decision result for an unauthorized state, the LLM collapses to the rejection response distribution $\delta_{rej}$ under the guidance of $T$.
5.2 Methodology
To achieve the above goals, the CoA framework systematically reconstructs LLMs’ information-processing methods along three dimensions: input semantic reorganization, output inference trajectory reconstruction, and training data synthesis and paradigm alignment. It aims to transform access control strategies from external defense mechanisms into internal cognitive capabilities.
Input restructuring and permission injection: To enable LLM to perceive heterogeneous permission boundaries, we restructured the input prompt structure. In the traditional paradigm, the model input only contains semantic content $I_{orig}=X\oplus E$, where $X$ is the user query and $E$ is the retrieved or tool calling feedback external context. System prompts and tool information are often treated as static background or external components, and LLMs lack information about identity and permissions.
Unlike existing methods, we incorporate permission information into the LLM input. By injecting user and tool permission tags into system-level prompts and adding explicit permission tags to the external context, we construct a clear permission context within the prompt words. The recombined input $I$ is defined as:
$X=Prompt_{sys}(C_{u},C_{g})\oplus Context(E,C_{e})\oplus X$ (8)
where $C_{u}$ and $C_{g}$ represent the permission tags for the user and available tools, respectively, and are dynamically injected into the system-level prompt as global constraints for inference; $Context(E,C_{e})$ then explicitly marks the permissions of external knowledge with the permission tag $C_{e}$.

### 03 Apr 2026 v2

@@ 2.2.2 Data Synthesis and Fine-tuning
We then apply supervised fine-tuning (SFT) on the synthesized data. By minimizing the standard negative log-likelihood loss, we strongly bind the authorization decision to the response generation:
@@ 3 Evaluation
In this section, by analyzing and presenting experimental results on Chain-of-Authorization (CoA), we discuss whether CoA can endow Large Language Models (LLMs) with robust intrinsic permission-control capabilities without compromising their performance on downstream tasks.
@@ 3.1 Model Utility on Authorized States
In this section, we evaluate the performance of the proposed Chain-of-Authorization (CoA) framework under authorization constraints across three different model architectures (Qwen3-1.7B, Llama3.1-8B-instruct, and Mistral-7B-v0.3). The evaluation benchmarks are divided into three categories: internal parametric knowledge tasks, including WMDP [li2024wmdp] and MMLU [hendryckstest2021mmlu]; external context tasks, including SQuAD [rajpurkar-etal-2016-squad] and CovidQA [covidqa, moller-etal-2020-covidqa]; and tool-calling tasks, represented by Mobile-Actions [google_mobile_actions]. We use accuracy to evaluate each model’s utility and policy adherence.
We compare CoA against several baselines, including the Base model, direct supervised fine-tuning (SFT), extra classification model (SFT+Extra), Permissioned LLM (PermLM), and sudoLM. The SFT+Extra method employs a RoBERTa-base classifier to identify the required permission in user prompts, so the fine-tuned LLM receives only relevant context in an authorized state. PermLM implements submodels for WMDP and MMLU and employs a context-filtering strategy in SQuAD and CovidQA to ensure LLMs receive contexts only from authorized documents.
@@ 3.2 Authorization under Different Permissions
Figure 2: Accuracy and rejection rate across five datasets under mismatch and public authorization states, where each circle represents the average accuracy of a method on a group of specific datasets across three backbone models, with the line mapping to its variance.
Figure 2 illustrates the accuracy of different methods in the unauthorized state. Since high accuracy in the unauthorized scenario usually means that unauthorized information has been successfully utilized, the ideal authorized method should result in the “Optimal Zone” in the lower left corner, that is, maintaining extremely low accuracy in both the public and mismatch states. Experimental results show that the Base and SFT models still maintain high accuracy in the unauthorized state, reflecting the lack of effective access control capabilities in the existing training paradigm. Although slightly lower than SFT in the public state, PermLM still retains high accuracy and does not completely block information leakage. The accuracy of sudoLM varies significantly across datasets and authorized states, indicating weak consistency. In contrast, the proposed CoA consistently reduces accuracy to near 0 across all datasets, with its results concentrated in the optimal zone, indicating its ability to effectively block the model’s access to knowledge and capabilities in the unauthorized state, thereby preventing unauthorized information leakage.
On the contrary, for the rejection rates shown in the lower 3 subplots, an ideal model should be located in the “Optimal Zone” in the upper-right corner, maintaining a high rejection rate in both the public and mismatch unauthorized states. As shown in Figure 2, the base model and SFT models lack basic authorization awareness, with an almost zero rejection rate across all datasets. While PermLM restricts information access through structural isolation, it also fails to establish an explicit rejection mechanism. The external module only demonstrates some rejection capability on tasks primarily based on parameterized knowledge (such as WMDP and MMLU), but its rejection rate drops significantly on tasks that depend on external context (such as SQuAD and CovidQA). Meanwhile, sudoLM’s performance is unstable, especially in the mismatch state, where it often continues to generate normal answers. In contrast, our proposed CoA consistently falls within the optimal zone on all models and datasets, achieving a rejection rate of nearly 100% in both the public and mismatch unauthorized states, demonstrating its ability to reliably identify fine-grained authorization states.
@@ 3.3 Robustness Against Adversarial Prompts
Table 2: Attack Success Rate (%) under the mismatch and public authorization state on WMDP. Bold values indicate the best results. Results are shown in the mismatch/public format.
@@
None 100.00/3.67 0.14/6.54 78.64/9.25 0.14/0.00 98.50/17.82 0.00/0.00
Prefix Injection 100.00/3.27 0.14/6.80 79.18/0.00 0.14/0.00 0.95/0.00 0.27/0.00
Style Injection 100.00/99.86 0.14/9.12 51.43/7.89 0.14/0.14 98.10/66.12 0.14/0.00
Misrepresentation 99.86/0.00 0.14/5.58 57.28/4.90 0.00/0.00 97.28/1.77 0.14/0.00
Logical Appeal 100.00/0.00 0.00/5.44 72.38/3.68 0.14/0.00 98.24/0.69 0.00/0.00
Authority Endorsement 100.00/0.00 0.00/5.31 66.12/1.50 0.00/0.00 97.42/0.41 0.00/0.00
Expert Endorsement 100.00/0.00 0.27/4.49 48.03/2.32 0.27/0.00 97.15/0.82 0.00/0.00
Evidence Persuasion 100.00/0.28 0.00/5.44 69.52/1.91 0.00/0.00 98.51/0.14 0.14/0.00
PAIR 6.39/0.82 0.00/0.68 0.27/0.27 0.00/0.27 12.52/2.99 0.00/0.00
As shown in Table 2, under the mismatch authorization state, CoA exhibited strong defensive consistency, obtaining the optimal or near-optimal ASR in most scenarios. Particularly in the style injection scenario, CoA successfully reduces the ASR of the Llama-3-8B-Instruct from $52.52\%$ to $0.14\%$. On the other hand, under the public authorization state, despite facing persuasive attacks driven by GPT-4o, CoA repeatedly achieved an ASR of $0.00\%$ on both the Llama-3.1-8B-Instruct and Mistral-7B-v0.3. In contrast, sudoLM showed significant fluctuations with style injection, with its ASR rising to $66.12\%$ on Mistral-7B-v0.3, while CoA remained robust at $0.00\%$, further highlighting its reliability in handling non-static authorization boundaries. In summary, this cross-model robustness demonstrates that CoA can effectively filter out semantic perturbations intended to circumvent authorization policies by embedding authorization decisions into the reasoning trajectory.
@@ 3.4 Visualization: What does CoA do?
To delve deeper into how the CoA mechanism reshapes the model’s internal representation space, we conducted a visualization analysis of the fine-tuned Llama-3.1-8B-Instruct model on the WMDP dataset, comparing it with its original version. First, we constructed examples under three authorization states using the test set:
@@
  • Public: We add the public permission label to all samples to simulate a visitor’s access scenario.
Subsequently, for the original model, we extracted the hidden state of the prompt’s last token in the last transformer layer. For the fine-tuned CoA model, we extracted features at two positions: one is the last token of the prompt, at which point the model has absorbed complete semantic and permission input information; the other is the last token before the authorization reasoning process ends and the final answer is generated, containing the authorization decision information after reasoning. Finally, we first use PCA to reduce the dimension to $50$ to suppress noise and improve computational efficiency [^1] We perform PCA first, following the official guidance of scikit-learn library https://scikit-learn.org/stable/modules/generated/sklearn.manifold.TSNE.html., and then use t-SNE to reduce the dimension to $2$ for visualization.
@@
The visualization results are shown in Figure 3. The original model (prompt end) shows the distribution of representations at the prompt’s end. Although the public samples form independent clusters, the match and mismatch samples highly overlap. This indicates that although the original model can recognize the permission-label format, it cannot semantically distinguish the prompt across different authorization states. CoA model (prompt end) shows the representation at the same position in the fine-tuned model. Representations at the end of the prompt still overlap, indicating that the model does not classify the authorization status based solely on permission labels in the input. In contrast, the CoA model (reasoning end) shows a completely different distribution of representations before the final response. The three sample types form distinct, non-overlapping clusters. This evolution from overlapping to separation demonstrates that CoA successfully decouples the authorization decision logic from the original semantics through the authorization reasoning process, and maps it to different representation regions, thereby enabling the model to accurately identify prompts under different authorization states.
@@ 3.5 Causal Analysis of CoA Trajectories
Figure 4 illustrates the evolution of the reasoning trajectory of CoA under targeted intervention and its causal impact on the final answer. The experiment injected conflict tokens into key stages of the authorization process, including the system prompt, resource review, identity resolution, and decision-making, and compared the model’s behavior under standard and intervention trajectories. The results show that under a normal baseline without intervention, the model can stably complete the task without rejection. However, when any key stage exhibits logical inconsistency (e.g., invalid resource or permission mismatch), the reasoning trajectory undergoes a systematic shift, ultimately triggering the rejection mechanism and increasing the rejection rate from $0\%$ to $100\%$. This result indicates that the security of CoA does not stem from superficial cue-word constraints but is driven by the causal dependencies among logical states in the authorization process. The internal consistency of the reasoning trajectory plays a crucial role in the authorization process, enabling the model to generate different responses depending on the authorization state.
@@ 3.6 Ablation Study
To systematically evaluate the impact of hyperparameters on model performance, we conducted experiments on the WMDP dataset using the Llama3-8B-Instruct model with batch sizes of $4$ and training for $5$ epochs. A cosine learning-rate scheduler was used, and the first $10\%$ of steps served as a linear warm-up phase to ensure optimization stability. We saved checkpoints at the end of each epoch and performed evaluation on the test set. The main evaluation metrics included accuracy (acc) and rejection rate (rej), covering three authorization states: match, mismatch, and public.
@@ 5 Methods
5.1 Permission-aware Data Synthesis
To internalize the aforementioned reasoning capabilities within the model, we designed a permission-aware data synthesis method. Based on the extent to which user permissions cover the permissions required for the model to generate compliant responses, we constructed three types of sample sequences $(C,X,T,Y)$ that cover different authorization states by permuting user permissions and model responses. This allows the model to learn permission-matching relationships:
  • Matched Authorization: In the matched authorization state, $C_{u}$ covers the permission of $E$ and satisfies the internal knowledge and permission of the tool to be called for prompt $X$. In this state, we use the union of the permission labels of prompts, external context, and the tool to be called as the user’s permission labels, i.e., $C_{u}=C_{e}\cup C_{x}\cup C_{g}$. The model responds normally to the user prompt.
<think>
The question is about [Prompt Permission] // Resource review stage
[Final Decision] // Decision-making stage
</think>
[Response] // Final output to user